OpenStudy (anonymous):

Ok this one is confusing me? Daughter needs help again. 7/8 divided by 1/12

OpenStudy (anonymous):

\[\dfrac{\dfrac{7}{8}}{\dfrac{1}{12}}=\dfrac{7}{8}\times\dfrac{12}{1}\]

OpenStudy (anonymous):

we got that far...its after that, that confuses me and I can't help her if I am confused.

OpenStudy (anonymous):

\[7\times 12 = 84\]

OpenStudy (christos):

you multiply 7 with 8 and 1 with 12

zepdrix (zepdrix):

Is the simplification the part that's confusing you?

OpenStudy (christos):

7 with 8 goes up and rest down

OpenStudy (anonymous):

Yes zepdrix

OpenStudy (anonymous):

\[8\times1=8\] therefore \[\dfrac{84}{8}=10.5\]

OpenStudy (anonymous):

Thank you Smart wish...need a little further explanation so hopefully I don't get confused again.

zepdrix (zepdrix):

\[\large \frac{84}{8}\]We're looking for factors that both numbers have in common.\[\large 4 \times 21=84\]\[\large 4 \times 2=8\] See how they both have a 4 in common? We can write our fraction this way,\[\large \frac{84}{8} \qquad = \qquad \frac{21\times4}{2\times 4}\]We broke down the 84 into 2 factors and id the same with the 8. Is that part confusing?

OpenStudy (anonymous):

That helps a little bit more breaking it down like that. I have a learning problem when it comes to math. LOL. I am trying.

zepdrix (zepdrix):

hehe :3

zepdrix (zepdrix):

\[\large \frac{21\times4}{2\times4}\] Whenever a value is divided by itself, it gives us 1. Example \(\large \dfrac{2}{2}=1\). The value 1 does not change our fraction in any way. So we can ignore it. So let's try that with the 4's. 4 divided by 4 gives us 1.\[\large \frac{21}{2}\times 1\]Since we're going to ignore the 1, it simplifies down to,\[\large \frac{21}{2}\]

zepdrix (zepdrix):

We sometimes refer to this process as "cancelling out".\[\large \frac{21\times \cancel4}{2\times \cancel 4}\]

OpenStudy (anonymous):

so I think i got the final answer 84/10= 8 4/10

zepdrix (zepdrix):

Oh the instructions say to write it as a mixed number?

OpenStudy (anonymous):

I know she has to change it from a improper to a mixed.

OpenStudy (anonymous):

Sorry should have specified.

zepdrix (zepdrix):

Oh ok :) Yes that answer looks good. But it appears we can simplify down the \(\large \dfrac{4}{10}\) a little bit.

OpenStudy (anonymous):

Yeah, i seen that too. so it would be 2/5

zepdrix (zepdrix):

\[\large 8\;\frac{2}{5}\] Something like that? Yay good job! :)

OpenStudy (anonymous):

that looks right. :) Gonna go on to the next one hopefully I don't get confused again. If I do I will be back. :)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

OMG....Okay if anyone is still paying attention to this question I am gonna go crazy. 3/4 divided by 9/16

OpenStudy (anonymous):

i was looking over the posts and you said you got 8 and 2/5, but before that it was said of 21/2 was gotten from simplifying. 84/8 = 21/2 = 10 and 1/2 so the 8 and 2/5 isnt correct

OpenStudy (anonymous):

and to the recent question (3/4) / (9/16) you want to make this just one number over another number. one thing in math is if you multiply by one you get the same number so also:

OpenStudy (anonymous):

then when you multiply fractions you multiply across so multiply numerators and denominators so 3 * 16 / 4 * 9 = 48 / 36 if you have trouble reducing fractions just start simple. you see they are both even so take a factor of 2 out of each 2 * 24 / 2 * 18 : so you took the 2 out of each and have 24/18 : still even so you can take another 2 out of each 2* 12 / 2 * 9 : take the 2 out again and have 12/9 : not even but still simple, go to the next number which is 3, and it works 3* 4/ 3* 3 : take out the 3 4/3